jasonmitchell:

Cydney is proving that perpendicular lines have slopes that are opposite reciprocals. She draws line p and labels two points on the line as (0,a) and (b,0) . Enter the answers, in simplest form, in the boxes to complete the proof. The slope of line p is ______. Rotate line p 90° counterclockwise about the origin to get line q. The labeled points on line p map to (−a,0) and (____, ____) on line q. The slope of line q is _____. The slopes of the lines are opposite reciprocals be

jasonmitchell:

@TheSmartOne

SmokeyBrown:

Well, the slope of line p would be a/b, since the y variable increases from 0 to a as the x variable increases from 0 to b. If you do the rotation as described, line q would have the point (0, b) on the y axis. And the slope of line q would be -b/a, by the same reasoning we used to find the slope of line p.

Cupcake123456:

Slope of p = -a/b (0, b) is on line q Slope of q = ( b - 0) / (0 - -a) = b/a Product of the slopes = (-a/b) (b/a) = (-ab) /(ab) = -1

Cupcake123456:

@SmokeyBrown u agree ?

jasonmitchell:

hmm convincing enough

SmokeyBrown:

Yeah, cupcake is right, I was wrong about the slopes

jasonmitchell:

nah both of you were good

jasonmitchell:

thx a lot man

Cupcake123456:

best responds ?

jasonmitchell:

yup